# ASTERIX: Module for modelling the water flow on vegetated hillslopes

### v1

@@ 4.2 Thacker Problem
Depending on the values of some parameters, the exact solutions of this problem can be classified into oscillating, non-oscillating and mixed ones. In [19], we have considered the case of Thacker Problem with an oscillating solution and compared it with the numerical one described there and implemented in ASTERIX. In this subsection, we present a comparison between a non-oscillating solution for Thacker Problem and the numerical solution provided by ASTERIX [^2] The soil resistance term $\mathfrak{t}^{s}_{a}$ is theoretically chosen to be proportional to the water depth and velocity in Thacker Problem in order to obtain an analytic solution. Note that this form (28) is different from (6), and ASTERIX was modified accordingly.. The non-oscillating solution was obtained using $a=1.25\cdot 10^{-3}$, $b=5\cdot 10^{-3}$, $x_{0}=y_{0}=500\ {\rm m}$ in (27), and a proportionality coefficient of the friction force $\tau=0.7\ {\rm s}^{-1}$ in (28). The reader is referred to B for extended mathematical details about the analytic solution to this problem.
@@
where $g=9.81\ {\rm m}\cdot{\rm s}^{-2}$ is the gravitational acceleration. Fig. 5 shows that our numerical solution (obtained on a regular hexagonal mesh with $487123$ cells of radius $0.888231\,{\rm m}$) is in accordance with the oscillating planar free water surface given by the analytic solution to the Thacker Problem.
$t=10\,{\rm s}$, $err=0.0107$ $t=30\,{\rm s}$, $err=0.0102$
@@
Figure 5: 2D Thacker Problem: dynamics of the numerical free water surface. The pictures were obtained using gnuplot on data calculated with ASTERIX. The error between the numerical and analytic solutions is $err=||((h+z)^{app}-(h+z)^{ex})/(h+z)^{ex}||_{\infty}$.
The evolution of the error $err=||((h+z)^{app}-(h+z)^{ex})/(h+z)^{ex}||_{\infty}$ between the exact and numerical solution is pictured in Fig. 6.
@@
Figure 6: 2D Thacker Problem: time evolution of the error $err=||((h+z)^{app}-(h+z)^{ex})/(h+z)^{ex}||_{\infty}$ of discharge between the analytic and numerical solutions for two different hexagonal cell sizes: $1.65\,{\rm m}$ and $0.888\,{\rm m}$ for the left and right picture, respectively.
4.3 Simulation on Şuşiţa River basin
@@
Unfortunately, we do not have data for the water distribution, plant cover density and measured velocity field in a hydrographic basin to compare our numerical results with. However, to be closer to reality, we have used the GIS data for an approximately $370\,{\rm km}^{2}$ soil surface included in a $64\times 28\,{\rm km}^{2}$ rectangular area. More precisely, this rectangular area, Fig. LABEL:fig_susita_foto, lies north-west of Tişiţa whose geographical coordinates are $45^{\circ}$ $50^{\prime}$ $36^{\prime\prime}$ North, $27^{\circ}$ $13^{\prime}$ $32^{\prime\prime}$ East.
We accomplished a theoretical experiment: starting with a uniform rainfall generated by a triangular model (see C for extended details) and characterized by
@@
on the entire basin and using two different uniform vegetation densities ($\theta_{1}=1.0$ and $\theta_{2}=0.97$), we ran ASTERIX on a network with $84651$ hexagonal cells of radius $41\,{\rm m}$.
The water distribution at a given moment in time is comparatively presented in Fig. 8 using a blue color gradient (the darkest blue corresponds to the cells with $\theta h\geq 4h_{0}$, where $h_{0}=i_{a}*T_{d}/(2*1000)=0.0366\,{\rm m}$ is the rainfall depth). The water velocity distribution at the same moment in time is pictured in Fig. 9 using a red color gradient for the modulus of the velocity and arrows for the flow direction. The convention of coloring the relief where $\theta h\leq 0.1h_{0}$ is applied for both figures. The data in Table 4 confirms that more water is retained in the basin if the vegetation is denser.
@@
$1.00$ $3006.6$ $23.8$ $2982.8$
$0.97$ $3006.6$ $20.7$ $2985.9$
Figure 8: Snapshots of water distribution on Şuşiţa’s hydrographic basin for two different uniform vegetation densities: $\theta=1.0$ and $\theta=0.97$ on the upper and bottom picture, respectively. As expected, our numerical data are consistent with terrain observations: the amount of water leaving the basin is greater in the case of lower vegetation density.
Figure 9: Snapshots of water velocity distribution on Şuşiţa’s hydrographic basin for two different uniform vegetation densities: $\theta=1.0$ and $\theta=0.97$ on the upper and bottom picture, respectively. As expected, the velocities are smaller when vegetation is denser.
@@
Remark 4.1.
@@
4.4 Sensitivity
Estimation of friction coefficients
The parameters $\alpha_{s}$ and $\alpha_{p}$ can not be directly measured, they must be estimated by other methods. Therefore, it is very important to know how its imprecise assessment affects the solution of the Saint-Venant equation.
To find these parameters, we use here an inverse method and the measurements of an experiment reported in [21]. The experimental installation consists of a rectangular laboratory flume ($18$ m long and $1$ m wide) with a longitudinal bottom slope $S=1.05$ mm/m and partially covered with vegetation. The vegetation is modelled using uniformly distributed emergent circular cylinders of radius $R_{c}=5\,{\rm mm}$ and has a density of $N=81\,{\rm cylinders}/{\rm m}^{2}$. This type of vegetation allows the estimation of the porosity by $\theta_{v}=1-N\pi R_{c}^{2}\approx 0.99364$. The experiment was performed for different given values of the upstream flow rate $q_{L}$, free downstream discharge, two types of vegetation distribution,
$\theta_{1}=\left\{\begin{array}[]{ll}1,&x\in[0,9)\\ \theta_{v},&x\in[9,18]\end{array}\right.,\quad\theta_{2}=\left\{\begin{array}[]{ll}\theta_{v},&x\in[0,9)\\ 1,&x\in[9,18]\end{array}\right.$
and $z=-m(x-18),\,x\in[0,18]$, with $m=0.105$ mm/m. The numerical simulation is carried out using ASTERIX with the following boundary conditions:
$uh(t,0)=q_{L},\quad\partial_{x}u(t,18)=\partial_{x}h(t,18)=0.$
We will perform later a sensitivity local analysis around of some reference values $(\alpha^{R}_{s},\alpha_{p}^{R})$ determined by fitting the experimental data reported in [21]. To illustrate the difficulties of estimating these friction parameters, we minimize the errors between the measured and calculated water depth on a $21\times 26$ grid inside the domain
$\Lambda:=\displaystyle\{(\alpha_{s},\alpha_{p})\big{|}\quad 0.0025\leq\alpha_{s}\leq 0.0185,\quad 55\leq\alpha_{p}\leq 80\}.$
In this sense, we define the error function
$\chi(\alpha_{s},\alpha_{p}):=\sqrt{\frac{1}{N_{e}N_{g}}\sum_{l=1}^{N_{e}}\sum_{i=1}^{N_{g}}\displaystyle\left(h^{\rm exp}(x_{i}^{l})-h(x_{i}^{l};\alpha_{s},\alpha_{p})\right)^{2}},$ (30)
$\chi(\alpha_{s},\alpha_{p})=\sqrt{\frac{1}{6\cdot 17}\sum_{j,k,i}\displaystyle\left(h^{\rm exp}(x_{i}^{k};q_{jk};\theta_{k})-h(x_{i}^{k};q_{jk};\theta_{k},\alpha_{s},\alpha_{p})\right)^{2}}.$ (31)
with
$j=\overline{1,3},\quad k=\overline{1,2},\quad i=\overline{1,17},$
and
@@ 4.4 Sensitivity
Fig. 11 presents the contour levels of the error surface for the following cases:
Case I: Experiment with $\theta=\theta_{1}$, $q=q_{21}$;
Case II: Experiments with $\theta=\theta_{1}$, $q\in\{q_{11},q_{21},q_{31}\}$;
Case I Case II
Plant friction: $\alpha_{p}$
Refer to caption Refer to caption
Case III Case IV
Plant friction: $\alpha_{p}$
Refer to caption Refer to caption
Soil friction: $\alpha_{s}$ Soil friction: $\alpha_{s}$
Figure 11: The contour level lines of the error surface in the four analyzed cases
When using only one experiment, e.g. Case I, the minimum is poorly located: there are many pairs $(\alpha_{s},\alpha_{p})$ in the area of minimum error and it is not possible to decide which one is the best. For Case II and Case III, each with 3 experiments, the parameters are not yet sufficiently well located: $\alpha_{s}$ is poorly located in Case II and $\alpha_{p}$ is poorly located in Case III. For Case IV, the surface of the error function (31) is drawn in Fig. 12 and the minimum is the better located.
Figure 12: The error surface for Case IV. The minimum of this function is reached at $\alpha_{s}=0.00709,\,\alpha_{p}=73.39$: $\quad{\rm min}=\chi\left(0.00709,73.39\right)=2.26$ mm.
For solving a minimization problem, more efficient methods can of course be used than searching for the solution on a grid. Since the evaluation of the cost function (31) associated with all six experiments at one single point requires a considerable computational effort, we used the “Optimization by pattern search” algorithm [34] based on the “Direct Search” method introduced in [35] and found
$\alpha_{s}=0.00709\quad{\rm and}\quad\alpha_{p}=73.39.$ (32)
A comparison between the measured data and numerical water depth found with these determined parameter values is illustrated in Fig. 13.
@@
For each experiment, the error $\epsilon=\chi\left(0.00709,73.39\right)$ (in [mm]) between the two solutions is calculated with (30) for $N_{e}=1$ and $N_{g}=17$; Table 5 presents the values of these errors.
@@
$q$ $q_{11}$ $q_{12}$ $q_{13}$ $q_{21}$ $q_{22}$ $q_{23}$
$\epsilon$ 1.48 2.05 1.73 3.08 2.10 2.73
@@
We further proceed and make a theoretical analysis for the response of the averaged water depth value along the vegetated zone $9\leq x\leq 18$ to the variation of the environmental variables $q$, $\theta_{v}$ and $m$. Let us introduce $\overline{h}\left((\alpha_{s},\alpha_{p});(q,\theta_{v},m)\right)$ by
@@
$\alpha^{i}_{s}=\underline{\alpha_{s}}+(i-1)\triangle_{s},\quad\alpha^{j}_{p}=\underline{\alpha_{p}}+(j-1)\triangle_{p},\quad i=\overline{1,N_{s}},\quad j=\overline{1,N_{p}}.$
The sensitivity is quantified by:
@@
Fig. 14 illustrates the behavior of the sensitivities $\delta_{s}$ (top row) and $\delta_{p}$ (bottom row) of the model with respect to the environmental variables $\boldsymbol{p}$. The porosity $\theta$ is fixed to $\theta_{v}=0.9936$ on the first column and the slope $m$ is fixed to the value $0.105$ on the second column. A first observation we can make is that, in general, the sensitivity increases with the discharge $q$. An exception can be noticed for $\delta_{s}$ at small values of $\theta_{v}$, see the picture B. As expected, a lower vegetation density leads to higher sensitivity with respect to the water-soil friction coefficient and lower sensitivity with respect to the water-plant friction coefficient, see pictures B and D. Also, as the longitudinal slope $m$ increases (the gravitational forces start to dominate), the sensitivity with respect to the water-plant friction coefficient diminishes, see picture C. We can also observe that the sensitivity with respect to the water-soil friction coefficient increases for small values of $m$ and decreases for higher values of $m$, see picture A.

### v2

@@ 4.1.2 Thacker’s Problem
Depending on the values of some parameters, the exact solutions of this problem can be classified into oscillating, non-oscillating and mixed ones. In [20], we have considered the case of Thacker’s Problem with an oscillating solution and compared it with the numerical one described there and implemented in ASTERIX. In this subsection, we present a comparison between a non-oscillating solution for Thacker’s Problem and the numerical solution provided by ASTERIX [^2] The soil resistance term $\mathfrak{t}^{s}_{a}$ is theoretically chosen to be proportional to the water depth and velocity in Thacker’s Problem in order to obtain an analytic solution. Note that this form (22) is different from (6), and ASTERIX was modified accordingly.. The non-oscillating solution was obtained using $a=1.25\cdot 10^{-3}$, $b=5\cdot 10^{-3}$, $x_{0}=y_{0}=500\ {\rm m}$ in (21), and a proportionality coefficient of the friction force $\tau=0.7\ {\rm s}^{-1}$ in (22). The reader is referred to Appendix B for extended mathematical details about the analytic solution to this problem.
@@
where $g=9.81\ {\rm m}\cdot{\rm s}^{-2}$ is the gravitational acceleration. Fig. 5 shows that our numerical solution (obtained on a regular hexagonal mesh with $487123$ cells of radius $0.888231\,{\rm m}$) is in accordance with the non-oscillating planar free water surface given by the analytic solution to the Thacker’s Problem.
@@
$t=70\,{\rm s},\qquad err=0.0122$ $t=330\,{\rm s},\qquad err=0.0045$
[…]
The evolution of the error
$err=||((h+z)^{app}-(h+z)^{ex})/(h+z)^{ex}||_{\infty}$
between the exact and numerical solution is pictured in Fig. 6.
@@
Figure 6: 2D Thacker’s Problem: time evolution of the error $err=||((h+z)^{app}-(h+z)^{ex})/(h+z)^{ex}||_{\infty}$ between the analytic and numerical solutions on two hexagonal networks with different cell sizes: $1.65\,{\rm m}$ and $0.888\,{\rm m}$ for the left and right picture, respectively.
Fig. 7 shows a comparison between the time evolution over time of the numerical and analytic solutions of the free water surface at the three different points $P_{1}$, $P_{2}$, $P_{3}$ from Table 5.
Table 5: Data for the three points in Tacker’s Problem where the evolution over time of the free water surface is drawn in Fig. 7.
$P_{1}$ $P_{2}$ $P_{3}$
$x$ $577.692$ $500.000$ $264.615$
$y$ $451.221$ $500.518$ $564.471$
$z$ $1.944$ 00 $9.004$
Figure 7: 2D Thacker’s Problem: time evolution of the free water surface at the three points from Table 5.
4.1.3 Test on a Riemann Problem with vegetation
Riemann Problem is a mathematical subject associated to the Dam Break Problem, a real physical phenomenon. It represents a very important issue in the theory of hyperbolic systems because one can find an analytical solution of the problem for a broad spectrum of initial data.
The analytic solution is set up by a bunch of very special solutions, namely, shock waves and simple waves. The combination of these waves in a packet is ruled by a physical admissibility criteria. In the absence of such criteria the Riemann Problem can have a multitude of solutions, some of them having no physical sense. A numerical scheme is considered to be at least physically acceptable if it correctly predicts both types of waves.
The existence of the solution and its very special structure make the Riemann Problem almost a mandatory test for any numerical scheme dedicated to the shallow water equation.
The objective of our test is twofold: one aim is to verify how well the numerical solution approximates the true solution of the problem, and the other one is to verify if the 2D numerical scheme furnishes a markedly 1D solution for appropriate initial data and boundary conditions.
In this section, we choose a very difficult problem where both the soil surface and the porosity function exhibit simultaneous jumps. The piecewise constant soil surface and porosity functions
$(\theta,z)(x,y)=\left\{\begin{array}[]{ll}(\theta_{L},z_{L}),&x\leq 9\\ (\theta_{R},z_{R}),&x>9\end{array}\right..$ (24)
are given on a $18$ m long and $1$ m wide horizontal rectangular channel. The piecewise constant initial data are given by
$\left.h(x,y)\right|_{t=0}=\left\{\begin{array}[]{ll}h_{L},&x\leq 9\\ h_{R},&x>9\end{array}\right.,$ (25)
$\left.\boldsymbol{v}(x,y)\right|_{t=0}=\left\{\begin{array}[]{ll}(u_{L},0),&x\leq 9\\ (u_{R},0),&x>9\end{array}\right.,$ (26)
whose numerical values are in Table 6.
Table 6: Data for the 2D Riemann Problem
$\theta_{L}$ $z_{L}$ $h_{L}$ $u_{L}$ $\theta_{R}$ $z_{R}$ $h_{R}$ $u_{R}$
- [m] [m] [m/s] - [m] [m] [m/s]
$0.8$ $1.0$ $0.2$ $5.00$ $1.0$ $1.2$ $0.6$ $1.33$
The exact solutions of Riemann Problem are difficult to find and have only been studied in the one-dimensional case. For the 2D test considered in this subsection, the problem of finding the solution can be reduced to a 1D case for which it was shown the existence and uniqueness, [40]. The water surface calculated with ASTERIX at $t=1.5\;{\rm s}$ and a comparison between the analytic solution and the numerical data on a longitudinal section are presented in Fig. 8.
Refer to caption Refer to caption
Figure 8: 2D Riemann’s Problem: free water surface found with ASTERIX at $t=1.5\;{\rm s}$ (left picture) and a comparison between the analytic and numerical solutions from a longitudinal channel section (right picture, where the vegetated and the bare parts of the soil are marked with green and white, respectively).
The value of the mean relative error
$\frac{1}{N}\sum_{i=1}^{N}\left|\frac{(h+z)_{i}^{app}-(h+z)_{i}^{ex}}{(h+z)_{i}^{ex}}\right|$
between the numerical (app) and exact (ex) values of the free water surface is 0.003785, where $N=1122$ is the number hexagonal centers along the longitudinal section.
We remark that the three shock waves developed by the exact solution are also caught by ASTERIX. Furthermore, one can observe that the hexagonal grid used by our software does not induce numerical adverse effects such as additional head losses due to the impossibility of the hexagons to align exactly with the channel walls.
4.2 Tests on real data
4.2.1 Simulation on Şuşiţa River basin with vegetation
@@
Unfortunately, we do not have data for the water distribution, plant cover density and measured velocity field in a hydrographic basin to compare our numerical results with. However, to be closer to reality, we have used the GIS data for an approximately $370\,{\rm km}^{2}$ soil surface included in a $64\times 28\,{\rm km}^{2}$ rectangular area. More precisely, this rectangular area, Fig. 9, lies north-west of Tişiţa whose geographical coordinates are $45^{\circ}$ $50^{\prime}$ $36^{\prime\prime}$ North, $27^{\circ}$ $13^{\prime}$ $32^{\prime\prime}$ East.
Figure 9: Şuşiţa’s hydrographic basin. Satellite image [41] and the terrain reconstructed (from GIS data) on a hexagonal network.
We accomplished a theoretical experiment: starting with a uniform rainfall generated by a triangular model (see Appendix C for extended details) and characterized by
@@
on the entire basin and using two different uniform vegetation densities ($\theta_{1}=1.0$ and $\theta_{2}=0.97$), we ran ASTERIX on a network with $82148$ internal hexagonal cells of radius $41\,{\rm m}$.
The water depth distribution at a given moment in time is comparatively presented in Fig. 10 using a blue color gradient (the darkest blue corresponds to the cells with $\theta h\geq 4h_{0}$, where $h_{0}=i_{a}*T_{d}/(2*1000)=0.0366\,{\rm m}$ is the rainfall depth). The water velocity distribution at the same moment in time is pictured in Fig. 11 using a red color gradient for the modulus of the velocity and arrows for the flow direction. The convention of coloring the relief where $\theta h\leq 0.1h_{0}$ is applied for both figures. The data in Table 7 confirms that more water is retained in the basin if the vegetation is denser.
@@
$1.00$ $13166903$ $104416$ $13062487$
$0.97$ $13166903$ $90808$ $13076095$
Figure 10: Snapshots of water depth distribution on Şuşiţa’s hydrographic basin at $t=1500\;{\rm s}$ for two different uniform vegetation densities: $\theta=1.0$ and $\theta=0.97$ on the upper and bottom picture, respectively. The first and the second column bar next to the figures are the color palettes used for the relief and water depth, respectively. As expected, our numerical data are consistent with terrain observations: the amount of water leaving the basin is greater in the case of lower vegetation density.
Figure 11: Snapshots of water velocity distribution on Şuşiţa’s hydrographic basin at $t=1500\;{\rm s}$ for two different uniform vegetation densities: $\theta=1.0$ and $\theta=0.97$ on the upper and bottom picture, respectively. The first and the second column bar next to the figures are the color palettes used for the relief and modulus of the water velocity, respectively. As expected, the velocities are smaller when vegetation is denser.
@@
Fig. 13 illustrates the amount of water leaving the basin over time for different mesh sizes. One can observe that the curves are getting closer to each other as the grid resolution increases. This suggests a convergence property of the numerical solution to the true solution of the model.
Remark 2.
@@
Figure 13: The evolution of the total water leaving the basin over time for different grid resolutions.
4.2.2 Flow over a slope with vegetation
The flow on vegetated areas is more complex than on bare soil, which may explain the limited availability of reported data. In this section, we compare the numerical results provided by ASTERIX with those reported in [22].
The experimental installation consists of a rectangular laboratory flume ($18$ m long and $1$ m wide) with a longitudinal bottom slope $S=1.05$ mm/m and partially covered with vegetation. The vegetation is modelled using uniformly distributed emergent circular cylinders of radius $R_{c}=5\,{\rm mm}$ and has a density of $N=81\,{\rm cylinders}/{\rm m}^{2}$. This type of vegetation allows the estimation of the porosity by $\theta_{v}=1-N\pi R_{c}^{2}\approx 0.99364$. Six experiments were performed for different given values of the upstream flow rate $q_{L}$, free downstream discharge, two types of vegetation distribution,
$\theta_{1}=\left\{\begin{array}[]{ll}1,&x\in[0,9)\\ \theta_{v},&x\in[9,18]\end{array}\right.,\quad\quad\theta_{2}=\left\{\begin{array}[]{ll}\theta_{v},&x\in[0,9)\\ 1,&x\in[9,18]\end{array}\right.,$
and $z=-m(x-18),\,x\in[0,18]$, with $m=0.105$ mm/m. The values of $q_{L}$ for the first three experiments where vegetation is present on the lower half of the channel ($\theta=\theta_{1}$) are
@@ 4.2.2 Flow over a slope with vegetation
while the ones for the last three experiments where vegetation is present on the upper half of the channel ($\theta=\theta_{2}$) are
@@
The numerical simulations are carried out using ASTERIX with the following boundary conditions
$uh(t,0)=q_{L},\quad\partial_{x}u(t,18)=\partial_{x}h(t,18)=0,$
and the following values of the friction parameters
$\alpha_{s}=0.00709\quad{\rm and}\quad\alpha_{p}=73.39.$
A comparison between the measured data and numerical water depth is illustrated in Fig. 14.
@@
For each experiment $l$, the error $\epsilon$ (in [mm]) between the two solutions is calculated with
$\epsilon:=\sqrt{\frac{1}{N_{g}}\sum_{i=1}^{N_{g}}\displaystyle\left(h^{\rm exp}(x_{i}^{l})-h(x_{i}^{l})\right)^{2}},\quad\quad l=\overline{1,6},$
where $N_{g}=17$ is the number of gauges placed inside the channel for measurements, $x_{i}^{l}$ is the position of the gauge $G_{i}$ on the $l^{\rm th}$ experiment, while $h^{\rm exp}$ and $h$ are the measured and calculated water depths of the stationary flow, respectively. Table 8 presents the values of these errors.
@@
$q_{L}$ $q_{11}$ $q_{12}$ $q_{13}$ $q_{21}$ $q_{22}$ $q_{23}$
$\epsilon$ [mm] 1.48 2.05 1.73 3.08 2.10 2.73
Fig. 15 illustrates the water depth and the velocity field on the entire channel for the last experiment from Table 8 where $\theta=\theta_{2}$ and $q_{L}=q_{23}$. The water depth presented here corresponds to the longitudinal profile found on the third row and second column in Fig. 14.
Refer to caption Refer to caption
Figure 15: Flow over a slope with vegetation: water depth [mm] and velocity field [m/s] obtained with ASTERIX for the sixth experiment where $\theta=\theta_{2}$ and $q_{L}=q_{23}$.
One can observe that the water depth (and velocity) longitudinal profile is the same for any longitudinal section along the channel. Thus, the hexagonal grid used by the software does not induce numerical adverse effects such as additional head losses due to the impossibility of the hexagons to align exactly with the channel walls.
4.3 Sensitivity analysis
@@
We proceed and make a theoretical analysis for the response of the averaged water depth value along the vegetated zone $9\leq x\leq 18$ to the variation of the environmental variables $q$, $\theta_{v}$ and $m$. Let us introduce $\overline{h}\left((\alpha_{s},\alpha_{p});(q,\theta_{v},m)\right)$ by
@@
$\alpha^{i}_{s}=\underline{\alpha_{s}}+(i-1)\triangle_{s},\quad\quad\alpha^{j}_{p}=\underline{\alpha_{p}}+(j-1)\triangle_{p},$
for $i=\overline{1,N_{s}}$ and $j=\overline{1,N_{p}}$. The sensitivity is quantified by:
@@
Fig. 16 illustrates the behavior of the sensitivities $\delta_{s}$ (top row) and $\delta_{p}$ (bottom row) of the model with respect to the environmental variables $\boldsymbol{p}$.
@@
The porosity $\theta$ is fixed to $\theta_{v}=0.9936$ on the first column and the slope $m$ is fixed to the value $0.105$ on the second column. A first observation we can make is that, in general, the sensitivity increases with the discharge $q$. An exception can be noticed for $\delta_{s}$ at small values of $\theta_{v}$, see the picture B. As expected, a lower vegetation density leads to higher sensitivity with respect to the water-soil friction coefficient and lower sensitivity with respect to the water-plant friction coefficient, see pictures B and D. Also, as the longitudinal slope $m$ increases (the gravitational forces start to dominate), the sensitivity with respect to the water-plant friction coefficient diminishes, see picture C. We can also observe that the sensitivity with respect to the water-soil friction coefficient increases for small values of $m$ and decreases for higher values of $m$, see picture A.
Estimation of friction coefficients
Parameter estimation of a physical model is of crucial importance to ensure the accuracy of the model. The choice of an algorithm for solving this problem depends on the model and the available observational data. For example, in [44], the authors propose an automatic calibration method to estimate some parameters in a fully distributed shallow water equation system. The method is based on the Gauss-Marquardt-Levenberg algorithm to minimize a cost function and assumes the evaluation of the first and second order derivatives of the objective function. Since the analytic evaluation is almost impossible, a discrete derivative evaluation must be developed. Here is the place where the sensitivity of the model with respect to parameters comes into play.
The parameters $\alpha_{s}$ and $\alpha_{p}$ can not be directly measured, they must be estimated by other methods. Therefore, it is very important to know how its imprecise assessment affects the solution of the Saint-Venant equation. To find these parameters, we use here an inverse method and the measurements of an experiment reported in [22].
To illustrate the difficulties of estimating these friction parameters, we minimize the errors between the measured and calculated water depth on a $21\times 26$ grid inside the domain
[…]
In this sense, we define the error function
$\chi(\alpha_{s},\alpha_{p}):=\sqrt{\frac{1}{N_{e}N_{g}}\sum_{l=1}^{N_{e}}\sum_{i=1}^{N_{g}}\displaystyle\left(h^{\rm exp}(x_{i}^{l})-h(x_{i}^{l};\alpha_{s},\alpha_{p})\right)^{2}},$
$\chi(\alpha_{s},\alpha_{p})=\sqrt{\frac{1}{102}\sum_{j,k,i}\displaystyle\left(h^{\rm exp}(x_{i}^{k};q_{jk};\theta_{k})-h(x_{i}^{k};q_{jk};\theta_{k},\alpha_{s},\alpha_{p})\right)^{2}}.$ (29)
with
$j=\overline{1,3},\qquad k=\overline{1,2},\qquad i=\overline{1,17},$
and
$q_{11}=7\,{\rm l/s},\qquad q_{21}=15\,{\rm l/s},\qquad q_{31}=21\,{\rm l/s},$
$q_{12}=7\,{\rm l/s},\qquad q_{22}=15\,{\rm l/s},\qquad q_{32}=50\,{\rm l/s}.$
For solving a minimization problem, we use the “Optimization by pattern search” algorithm [45] based on the “Direct Search” method introduced in [46]. We consider the following cases:
Case I: Experiment with $\theta=\theta_{1}$, $q=q_{21}$;
Case II: Experiments with $\theta=\theta_{1}$, $q\in\{q_{11},q_{21},q_{31}\}$;
When using only one experiment, e.g. Case I, the minimum is poorly located: there are many pairs $(\alpha_{s},\alpha_{p})$ in the area of minimum error and it is not possible to decide which one is the best. For Case II and Case III, each with 3 experiments, the parameters are not yet sufficiently well located: $\alpha_{s}$ is poorly located in Case II and $\alpha_{p}$ is poorly located in Case III. For Case IV, the surface of the error function (29) is drawn in Fig. 17 and the minimum is the better located:
$\alpha_{s}=0.00709\quad{\rm and}\quad\alpha_{p}=73.39.$ (30)
Figure 17: The error surface for Case IV. The minimum of this function is reached at $\alpha_{s}=0.00709,\,\alpha_{p}=73.39$:
${\rm min}=\chi\left(0.00709,73.39\right)=2.26$ mm.